OpenStudy (anonymous):

Solve for x: -2(x + 3) = 4(x - 6)

mathslover (mathslover):

fist try to calculate -2(x+3) ..

mathslover (mathslover):

can u do this ? -2(x+3)=?

mathslover (mathslover):

always remember : a(b+c)=ab+ac

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

-2[x=3]=4[x-6] -2x-6=4x-24 2x=18 x=9

mathslover (mathslover):

@moha_10 please no direct answers @HillDP good answer bet please let the user do it him/her self

OpenStudy (anonymous):

that's true and im a girl :)

mathslover (mathslover):

k !

mathslover (mathslover):

so do u know what is -2(x+3) ?